solve for x
4x(x-8)-5x(x-7)=-4

Answer by checkley77(12844)

4x(x-8)-5x(x-7)=-4
4X^2-32X-5X^2+35X+4=0
-X^2+3X+4=0
X^2-3X-4=0
(X-4)(X+1)=0
X-4=0
X=4 ANS.
X+1=0
X=-1 ANS.